In millikan experiments and Oil Drop of radius10-4cm remain suspended between the plates which are 1cm apart.i f the drophas charge over it calculate the potential difference between the plates .the density of oil may be taken as 1.5gcm-3.

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

The amount of charge over the drop is not given. However forqamount of charge, the electric force on the drop must be equal to its weight.
Electric field between the plates=E=\dfrac{V}{d}

Hence electric force on drop=qE=q\dfrac{V}{d}=mg=\rho(\dfrac{4}{3}\pi r^3)g
\implies V=\dfrac{4\pi \rho r^3 gd]}{3q}